OpenStudy (anonymous):

Evaluate the Definite Integral:

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/2} (\sin 2Pit)/T - \alpha dt\]

OpenStudy (anonymous):

Please help! :)

OpenStudy (experimentx):

ask wolframalpha ... first

mathslover (mathslover):

\[\int\limits_{0}^{\pi/2} (\sin 2 \pi t)/T - \alpha dt\] *replace pi with the symbol

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i= \int\limits_{0}^{\pi%2F2}+%28\sin+2Pit%29%2FT+-+\alpha+dt&dataset=

mathslover (mathslover):

More modification : \[\large{\int\limits_{0}^{\pi/2} \frac{\sin 2 \pi t}{T} - \alpha dt}\]

OpenStudy (turingtest):

\[T=t\]?

OpenStudy (anonymous):

yup, the modified equation is correct.

OpenStudy (turingtest):

so big T is different than little t, right?

OpenStudy (anonymous):

Yup

mathslover (mathslover):

\[\large{T\ne t \space \textbf{or}\space t = T}\] ok so it is first one T is not equal to t

OpenStudy (anonymous):

Correct

OpenStudy (turingtest):

so treat big T as a constant

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Then U = 2PIt/T - Alpha, therefore, du = 2pi/T - alpha?

OpenStudy (turingtest):

\[\int_0^{\pi/2}\frac{\sin (2 \pi t)}{T} - \alpha dt=\frac1T\int_0^{\pi/2}\sin(2\pi t)dt-\alpha \int_0^{\pi/2}dt\]

OpenStudy (turingtest):

for the first integral\[u=2\pi t\implies du=2\pi dt\]the next should be pretty straight forward

OpenStudy (anonymous):

Okay, thank you! :)

OpenStudy (turingtest):

welcome!

Parth (parthkohli):

I love Substitution Rule! :)

OpenStudy (anonymous):

Can I solve it and double check what I got with you?

Parth (parthkohli):

... or do I hate it?

OpenStudy (anonymous):

I know, I hate it. lol.

OpenStudy (turingtest):

go ahead and try to solve it

OpenStudy (anonymous):

@TuringTest is the answer 1?

OpenStudy (turingtest):

no, that is not possible... let me see what it is (will be an ugly number)

OpenStudy (anonymous):

Oh Okay :(.

OpenStudy (turingtest):

yeah, ugly answer... first do\[\frac1T\int_0^{\pi/2}\sin(2\pi t)dt\]what do you get?

OpenStudy (anonymous):

I can't view the equation for some reason. I'll keep refreshing the page.

OpenStudy (anonymous):

Nope, still can't view it. I can only see the codes.

OpenStudy (anonymous):

@TuringTest I got -0.159 for that one.

OpenStudy (anonymous):

Sorry for the late reply. My net is acting up.

OpenStudy (anonymous):

Where does 1/T go ?

OpenStudy (anonymous):

Oh wait, so it's -1/T 0.159

OpenStudy (anonymous):

I'm sorry, I'm so bad at this.

OpenStudy (anonymous):

Can you ∫ sin 2π tdt ?

OpenStudy (anonymous):

Yeah, I got (-cos2π t)/2π t

OpenStudy (anonymous):

(-cos2π(π/2))/2π(π/2) - ((-cos2π)(0))/2π(0)

OpenStudy (anonymous):

Is that correct?

OpenStudy (anonymous):

u = -cos2πt, du = sin2π, 1/2du = sinπ?

OpenStudy (anonymous):

sorry about that.

OpenStudy (anonymous):

I'm not sure what to do after, definite integrals make no sense to me, again, I'm sorry.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yes, I can.

OpenStudy (anonymous):

u = 2π correct?

OpenStudy (anonymous):

2πt*

OpenStudy (anonymous):

so du = 2π = 1/2 du = π

OpenStudy (anonymous):

From TuringTest's instruction: u= 2πt -> du=2πdt Since you want dt: dt = du/2π

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Now plug into the equation (1/T )∫ sin 2π tdt = ... ?

OpenStudy (anonymous):

-1/T (cos 2πt)/2πt?

OpenStudy (anonymous):

Actually = (- 1/2πT ) ( cos 2πt )

OpenStudy (anonymous):

Okay, so this is when I plug in π/2 and 0?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

1.418T?

OpenStudy (anonymous):

I have class in 10 minutes, so I better go.

OpenStudy (anonymous):

I appreciate the help!

OpenStudy (anonymous):

Okay. Thanks again! :)